On January 1, 2021, Rapid Airlines issued $205 million of its 6% bonds for $188 million. The bonds were priced to yield 8%. Interest is payable semiannually on June 30 and December 31. Rapid Airlines records interest at the effective rate and elected the option to report these bonds at their fair value. On December 31, 2021, the fair value of the bonds was $194 million as determined by their market value in the over-the-counter market. Rapid determined that $1,000,000 of the increase in fair value was due to a decline in general interest rates.

Required:

1. to 3. Prepare the journal entries to record interest on June 30, 2021 (the first interest payment), on December 31, 2021 (the second interest payment) and to adjust the bonds to their fair value for presentation in the December 31, 2021, balance sheet. (Enter your answers in whole dollars. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

Answer 1 to 3:

General Journal Debit Credit No Date June. 30, 2021 Interest Expense $7,520,000 Discount on bonds payable $1,370,000 $6,150,0

Working:

1. June 30, 2021

Interest expense = 188000000 * 8%/2 = $7,520,000

Cash paid = 205000000 * 6%/2 = $6,150,000

Discount on bond payable = 7520000 - 6150000 = $1,370,000

Carrying value of bonds = 188000000 + 1370000 = $189,370,000

2. Dec 31, 2021

Interest expense = 189370000 * 8%/2 = $7,574,800

Cash paid = 205000000 * 6%/2 = $6,150,000

Discount on bond payable = 7574800 - 6150000 = $1,424,800

Carrying value of bonds = 189370000 + 1424800 = $190,794,800

3. Dec 31, 2021

Fair value adjustment = 194000000 - 190794800 = $3,205,200

Unrealized holding loss - NI = $1,000,000

Unrealized holding loss - OCI = $3,205,200 - $1,000,000 = $2,205,200